If a-15=c and a-5=b and b=?  then value of a+b+c=122
 
 
 

Grade:6

1 Answers

Khimraj
5 years ago
a -5 = b
so a = 5 + b    – (1)
a – 15 = cc
so c = (5+b) – 15 = b – 10       – (2)
and a + b + c = 122           – (3)
from equation (1), (2) and (3)
(5+b) + b + (b -10) = 122
3b = 127
so b = 127/3
Hope it clears.